OLIVER69:

find the x-intercepts from these equations. -4x^2 - 3x +1 = 0 2x^2 - 4x - 8 = 0

jhonyy9:

so like first step in case of the first equation divide both sides by -1 and will get 4x^2 +3x -1 = 0 in case of second equation factorize out the 2 2(x^2 -2x -4) = 0 divide both sides by 2 x^2 -2x -4 = 0